This a logbook of all the actions taken by the PH-DT group at the GIF

Updated by Sander Rouwette, Ivan Glushkov, Roberto Guida, Beatrice Mandelli

GIF RPC logbook

07.07.2011 - Friday

Time Creator Event
09:00 BM started a new set of measurements

02.07.2011 - Sunday

Time Creator Event
18:00 BM, RG P1, P2, P3 conditioning started

14.06.2011 - Tuesday

Time Creator Event
15:00 BM purifiers removed

12.06.2011 - Saturday

Time Creator Event
18:00 BM, RG switch off RPC 3up

30.05.2011 - Monday

Time Creator Event
12:00 BM regeneration of GC columns

11.05.2011 - Wednesday

Time Creator Event
18:00 BM, RG switch to closed loop mode the four new gaps

26.04.2011 - Tuesday

Time Creator Event
19:30 BM, RG hv restarted after gas kill (due to windows update)

11.04.2011 - Monday

Time Creator Event
10:00 BM, RG unexpected powercut
16:20 BM, RG power restarted after intervention from electrical piquet
16:30 BM, RG source ON, HV ON

21.03.2011 - Monday

Time Creator Event
18:50 BM, RG hv ON, source ON

18.03.2011 - Friday

Time Creator Event
11:30 BM, RG Normal mixture ON
11:30 BM, RG humidifiers refilled

11.03.2011 - Friday

Time Creator Event
19:50 BM, RG Argon ON

10.03.2011 - Thursday

Time Creator Event
09:30 BM hv off (pc reboot for unknown reason), gas ok. hv restarted

09.03.2011 - Wednesday

Time Creator Event
12:00 BM, RG humidifiers refilled

07.03.2011 - Monday

Time Creator Event
18:30 BM, RG humidifiers refilled

06.03.2011 - Sunday

Time Creator Event
22:30 BM gas pc restarted, HV restarted

04.03.2011 - Friday

Time Creator Event
14:30 BM humidifiers refilled

03.03.2011 - Thursday

Time Creator Event
9:30 BM hv restarted after gas kill

01.03.2011 - Tuesday

Time Creator Event
10:30 BM humidifiers refilled
11:00 BM weather station restarted

28.02.2011 - Monday

Time Creator Event
8:30 RG hv off (pc reboot), gas ok. hv restarted

25.02.2011 - Friday

Time Creator Event
9:15 BM found hv off, gas ok. hv restarted
11:00 BM hv restarted after gas kill (problem due to reboot of windows pc for unknown reason where the gas monitoring software is running)

22.02.2011 - Tuesday

Time Creator Event
12:00 BM,RG hv restarted after gas kill (problem due to automatic update of windows pc where the gas monitoring software is running)
15:30 BM,RG hv restarted after gas kill (problem due to some unannounced maintenance on the N2 dewar)
15:30 BM,RG filled fresh mixture humidifier
15:30 BM,RG filled tent humidifiers
19:00 BM,RG hv restarted after gas kill (problem due to automatic update of windows pc where the gas monitoring software is running ... second update of the day)

21.02.2011 - Monday

Time Creator Event
15:00 BM,RG new purifier configuration installed: P1(5A,4A),P2(4A),P3(R11)

18.02.2011 - Friday

Time Creator Event
16:30 RG humidifiers refilled
19:30 RG P1,P2,P3 conditioning started

13.02.2011 - Sunday

Time Creator Event
20:30 RG, BM humidifiers refilled

02.02.2011 - Wednesday

Time Creator Event
20:30 RG, BM GIF restarted after CERN general power cut

13.01.2011 - Thursday

Time Creator Event
20:30 RG resistivity measurement (rho24) ok for old gap

17.12.2010 - Friday

Time Creator Event
20:30 RG resistivity measurement (rho23)
21:40 RG gif off

14.12.2010 - Tuesday

Time Creator Event
20:00 RG hv kill and mixture monitoring reestablished.
  RG weather station still not working

13.12.2010 - Monday

Time Creator Event
10:00 RG PC used for kill and mixture monitoring: expected intervention from helpdesk today.

12.12.2010 - Sunday

Time Creator Event
21:30 RG GIF and gas on. PC used for kill and mixture monitoring is not working anymore. Expected intervention from helpdesk today.

11.12.2010 - Saturday

Time Creator Event
7:30 RG GIF and gas off for cern power cuts expected during all day. PC off and disconnected from 220 V

10.12.2010 - Friday

Time Creator Event
9:00 RG found HV off (not due to gas kill). HV restarted at 11 am
18:30 RG humidifier filled

09.12.2010 - Thursday

Time Creator Event
07:00 RG CERN general power cut
08:00 RG all restarted

07.12.2010 - Tuesday

Time Creator Event
09:00 RG humidifier filled

06.12.2010 - Monday

Time Creator Event
09:00 RG humidifier filled

03.12.2010 - Friday

Time Creator Event
18:30 RG humidifier repaired and filled

02.12.2010 - Thursday

Time Creator Event
15:30 RG problem with humidifier (broken)

01.12.2010 - Wednesday

Time Creator Event
3:00 RG gas kill. iC4H10 down to 1%, problem with pressure reducer. fixed at 11:30. HV restarted at 3:30 pm
15:30 RG gap1up trip at 1.5 kV. Cable ok. ... HV off
18:20 RG gaps GT-344 and GT-345: HV on after the gas KIll of this night, HV=9KV, I=33, 5.1uA

30.11.2010 - Tuesday

Time Creator Event
6:00 RG gas kill. However, at 8am the gas was flowing and mixture correct. HV restarted immediatly.
15:30 RG humidifier filled

--+++ 29.12.2010 - Monday

Time Creator Event
18:30 RG humidifier filled

27.11.2010 - Saturday

Time Creator Event
20:00 RG humidifier filled. T & RH adjusted

27.10.2010 - Wednesday

Time Creator Event
8:40 SR Door of the GIF was forced which resulted in a stop for 1.5 day.

20.10.2010 - Wednesday

Time Creator Event
13:00 SR Turned on the voltage something turned them off.
14:00 SR Changed the settings of the air conditioning and turned on the humidifier.

05.10.2010 - Wednesday

Time Creator Event
8:30 SR MDT started to install their chambers result in half our currents through the chambers
8:30 SR Started to notice that when the GIF control computer is stuck the sms service does not works.

01.09.2010 - Wednesday

Time Creator Event
14:41 RG & EA HV turn-off problem (seems to be) solved. Channels were not turned off around 3:30pm

27.08.2010 - Friday

Time Creator Event
16:00 RG & EA GC pump joint was greased.

25.08.2010 - Tuesday

Time Creator Event
16:00 RG & SR & EA Changed the filaments of the GC. Tried to see if the problem with the 2 dw gap was caused by the hv connectors

19.08.2010 - Thursday

Time Creator Event
9:00 SR GIF PC restarted

18.08.2010 - Wednesday

Time Creator Event
9:00 SR The filament of the GC broke.
11:30 SR Gif computer restart
15:30 SR Exactly at 15:30 the voltages went off again. Tommorow I will just before this time unplug the lan cable.

17.08.2010 - Tuesday

Time Creator Event
9:00 SR Restart of the computer
11:00 SR Restart of the computer
15:30 SR The currents went off again at 15:30.
15:35 SR After starting the HV again it went off again.
15:40 SR Currents are on a normal value again.
    If tommorow the currents will be turned off at approximatly the same time I will unplug the Lan Cable for 1 day, to see whether to problem presists or not.

16.08.2010 - Monday

Time Creator Event
9:00 SR HV's are still being turned off. Will start a closer monitoring. turned them on at 9:20 again.
9:00 SR GIF computer restarted.
10:30 SR The air in the system comes from the C2H2F4.
15:30 SR Currents approximatlty went off at 15:30. Turned them back on.

11.08.2010 - Wednesday

Time Creator Event
9:00 sr Found the HV off again. Problem with the power suply still exists.
9:30 SR Analysing the freshmix sample shows a 260x higher concentration of air inside the mixture than normal. Turned HV to a lower level

21.07.2010 - Wednesday

Time Creator Event
16:00 SR Power back on. Started the gas system again.
16:30 SR Started the GIF.

19.07.2010 - Monday

Time Creator Event
9:30 SR CLRETURN is fine at the moment.
11:30 SR Total power shut down. Everything is off including the GIF and it's control room

05.07.2010 - Monday

Time Creator Event
8:30 SR Voltages back up

21.06.2010 - Monday

Time Creator Event
9:00 SR GIF PC restarted

16.06.2010 - Wednesday

Time Creator Event
8:30 SR Voltages back up

15.06.2010 - Tuesday

Time Creator Event
9:00 SR The source has been off since 9:00 until 15:15. People from ATLAS were doing tests. Due to the fact that they will change to different absorbtion factors I lowered the voltages.

14.06.2010 - Monday

Time Creator Event
12:00 SR MS saturated
21:00 RG GIF PC need to restarted, but is not possible to access Bldg 190 (X-ray test).
    tomorrow we will start performing the HV scan with Source off.

03.06.2010 - Monday

Time Creator Event
10:30 SR GIF PC restarted.

03.06.2010 - Wednesday

Time Creator Event
9:00 SR GIF PC restarted.
14:00 RG & SR Replaced the He bottle for the GC.
20:10 RG GIF PC restarted.

01.06.2010 - Tuesday

Time Creator Event
9:00 SR GIF PC restarted.
20:00 RG GIF PC restarted.

27.05.2010 - Thursday

Time Creator Event
9:30 RG glass RPC at nominal high voltage.
17:30 RG Source OFF
17:45 RG Gas OFF for bakelite RPC test at nominal hv.
19:30 RG Gas ON for bakelite RPC test at nominal hv.
20:00 RG radiation ON
20:45 RG glass RPC test at 5 kV.

26.05.2010 - Wednesday

Time Creator Event
11:30 SR GIF PC restarted.
18:30 RG bakelite RPC at 6.0 kV Source OFF
18:45 RG Gas OFF for bakelite RPC test at 6kV.
20:20 RG Gas ON for bakelite RPC test at 6kV.
21:00 RG radiation ON
21:40 RG bakelite RPC test at nominal high voltage.
21:40 RG glass RPC test at 5 kV.

GIF RPC logbook

24.05.2010 - Monday

Time Creator Event
11:00 RG GIF PC restarted.
  RG bakelite RPC at 7.0 kV
16:00 RG Gas OFF for bakelite RPC test at 7kV.
17:30 RG Gas ON for bakelite RPC test at 7kV.
20:30 RG radiation ON
21:45 RG bakelite (and glass) RPC test at nominal high voltage.

23.05.2010 - Sunday

Time Creator Event
14:00 RG GIF PC restarted.
  RG Radiation OFF for several tests until tomorrow evening:
RG 1) long HV scan Glass RPC
RG 2) long run ATLAS-MDT
RG 3) RPC "mapping" open-closed-hv maps
  RG HV OFF for bakelite RPC until tomorrow for further tests.
23:00 RG GIF PC restarted.

21.05.2010 - Friday

Time Creator Event
16:30 RG GIF PC restarted.

19.05.2010 - Wednesday

Time Creator Event
11:30 SR GIF PC restarted.
16:30 SR PICO crashed restarded

18.05.2010 - Tuesday

Time Creator Event
09:30 SR GIF PC restarted.
12:00 SR GIF PC restarted.
14:30 RG Start test at 7 kV without gas flow
16:30 RG End test at 7 kV without gas flow. Restart gas flow
20:30 RG HV back to normal values

17.05.2010 - Monday

Time Creator Event
08:30 SR GIF PC restarted.

13.05.2010 - Thursday

Time Creator Event
14:30 RG GIF PC restarted.

11.05.2010 -Tuesday

Time Creator Event
13:30 SR GC PC crashed, did a reboot of the pc.
14:00 SR Pico PC was stuck, rebooted the pc.  

09.05.2010 - Sunday

Time Creator Event
17:30 RG GIF PC restarted.

08.05.2010 - Saturday

Time Creator Event
01:15 IG Gas system was in off state - restarted. BINOS always showing 5.4% - adjusted. Slow Conrtol PC stuck - restarted. chambers' HV - on
17:00 RG BINOS 5.0%

07.05.2010 - Friday

Time Creator Event
8:10 IG GIF PC restarted. GC analysis started.

06.05.2010 - Thursday

Time Creator Event
17:30 SR Entered the GIF to refill the humidifier

04.05.2010 - Tuesday

Time Creator Event
8:30 IG GIF PC restarted. GC analysis started. GC was in on state over the night.
10:30 SR, IG GIF PC restarted.

03.05.2010 - Monday

Time Creator Event
10:00 IG, SR GIF PC restarted. CAEN restarted.

30.04.2010 - Friday

Time Creator Event
10:45 SR GIF PC restarted.
15:00 RG, IG changed bubbler for glass rpc

29.04.2010 - Thursday

Time Creator Event
10:30 SR GIF PC restarted.
12:00 SR,RG HV scan radiation off
16:20 SR,RG HV scan ABS5
19:20 RG HV scan ABS2
17:20 IG GIF PC restarted.
24:00 RG HV scan ABS1

28.04.2010 - Wednsday

Time Creator Event
9:30 IG GIF PC restarted by Sander
15:00 SR Stopped and started the GIF to get the daqrpc reading right again. It was 0 for the source status while the source was on.

27.04.2010 - Tuesday

Time Creator Event
11:00 RG glass RPC back to nominal HV
12:00 IG+SR Refill of the humidifier, lower the dehumidifier setting, and Set the AC to 21 degrees.
  IG Connected the purifiers in OM in series. The first analysis will be done today.
  IG The manual integration is already possible on the laptop of Sander. To get all the automatic integration buttons active, all the methods from the GC laptop were copied locally.
19:30 IG GIF PC restarted.

26.04.2010 - Monday

Time Creator Event
07:45 IG GIF pc restarted. pcta1gs27 also restarted (was non-responsive). Started FRESHMIX analysis for Sander.
18:00 RG Installed bubbler on glass RPC return.
  RG Starting from today, do not take GC analysis on glass RPC return

25.04.2010 - Sunday

Time Creator Event
16:30 RG GIF pc restarted

23.04.2010 - Friday

Time Creator Event
9:30 IG+SR 7up in trip - restarted.
  IG+SR The temperature was 17deg (setting on the AC - 20 deg). The setting on the AC was changed to 23 deg.
  IG+SR The humidity was varying a lot (48-67%). We test a changed position of the dehumidifier (on top of the wall) in order to get access to more air.
11:00 IG SRA claims that the MS high-pressure pump has serious problems. Continuing only with the MS. What do we do in case of new impurities?.
12:00 IG Bought new external backup HDD for the MS/GC and a weather station. Both - to be installed.
17:30 IG GIF stopped for MDT calibration run.
18:00 IG OM cartridges installed for conditioning with 7 l/h

22.04.2010 - Thursday

Time Creator Event
8:30 RG bakelite gaps back to nominal HV
  RG glass gaps @ 7 kV at least until midday
15:15 IG 4up in trip, retsrted; 6up/down - overcurrent, adjusted limits to 100uA. Both are being observed.
15:40 IG+SR The temperature was high (23.5) and the RH was low (35%) Thus we switched off the heating and switched on the air condition, humidifier and dehumidifier.
16:30 IG Due to construction works near the GIF, the HV for the whole building was off and there was a fire alarm too.
    The MS/GC was properly stopped, and then when the power went back on - properly started.
    The distribution valves did not start again. Putting manually the offset 0.3 helped. The system is back to normal again, and starting to switch on the HV slowly.
20:30 IG Reset of the GIF PC.
23:00 RG all gaps back to normal HV

21.04.2010 - Wednesday

Time Creator Event
08:00 RG Found HV off for all gaps. No clear reason. restarting HV slowly.
09:45 RG All gaps back to nominal hv
11:30 IG Gas alarm ~ 10:00. I called Andre Castell, Franck Schmitt, David Hay and Miguel Riesco to get the problem fixed. What was found is that the alarm is strange - there are blinking signs "Gas Alarm", but there is no syren. There is no detection head activated, but the source was switched off. The gas people are still investigating. The GIF stays off.
12:00 all problem in restarting the system. Negative chamber pressure is producing an interlock (backup gas is ok)
12:30 IG + SR The heating in the tent is switched off, the humidifier and the dehumidifier are switched on.
15:00 all problem with interlock solved (offset added only for the start-up). Problem still need some investigation
17:30 all an interlock is produced by the HighVoltage and Gas monitoring software, but, looking at the MFCs controller everything is fine (no reson for the interlock)
19:30 RG also the total flow displayed by the HVC software is wrong since lunch time
20:00 RG apparently the ADC boot-up in some wrong status after the power-cut. After a reboot of the pc the problem was fixed
  RG pressure for return of open-mode is very unstable. After some tuning it is fine
21:00 RG Everything look fine. start to restart HV slowly.

16.04.2010 - Friday

Time Creator Event
09:00 IG+SR Restart of the GIF PC
10:32 IG+SR The supply bottle of the C2H2F4 was obviously changed - increase of C2H3F3, CH2F2 and Air - that means column B, retention times 18,9 - 32,42 - 45,20
12:00| IG+RG| The glass RPCs were switched on on steps - 5-6-6.5-7-7.4 kV. The currents are ok. It is not clear why they were high.|

14.04.2010 - Tuesday

Time Creator Event
09:00 RG I’m trying to switch on the glass rpc very slowly. So, if you see HV lower than 7.k kV, do not increase to 7.4 kV.
09:50 IG GIF PC restarted.

13.04.2010 - Tuesday

Time Creator Event
07:50 IG GIF PC restarted.
10:45 IG 7up tripped. Restarted. The current is high again! The same as 7down, ~7uA.
14:15 IG 7dw tripped. Restarted.
15:02 IG 7dw tripped again. Both channels show increase of current (from 6uA to 12uA and from 4uA to 12uA).

12.04.2010 - Monday

Time Creator Event
07:35 RG GIF PC restarted.
11:45 SR GIF PC restarted.
18:10 IG GIF PC restarted. Even the GIF logbook was not working - the entries were not visible on the screen.

11.04.2010 - Sunday

Time Creator Event
11:15 RG GIF PC restarted.
  RG found gap 3up in trip. ITrip was set to low. HV re-started

07.04.2010 - Wednesday

Time Creator Event
08:45 SR GIF PC restarted.

06.04.2010 - Tuesday

Time Creator Event
08:30 IG GIF PC restarted.

05.04.2010 - Monday

Time Creator Event
12:30 RG Found gas system off. Restarted. HV off until this evening or tomorrow
22:00 RG HV on

01.04.2010 - Thursday

Time Creator Event
11:30 RG HV scan vs ABS 1

31.03.2010 - Wednsday

Time Creator Event
11:10 All Change of the MS purifier in the closed loop.
  IG+SR Switch on the heater and the humidifier.
  IG Flow through the test purifier changed from 5 to 20 l/h.
21:50 RG Restart GIF Control PC

30.03.2010 - Tuesday

Time Creator Event
- SR+RG HV scan vs ABS 2,5,off

27.03.2010 - Saturday

Time Creator Event
23:50 RG Restart GIF Control PC

25.03.2010 - Thursday

Time Creator Event
9:30 IG Restart GIF Control PC

24.03.2010 - Wednesday

Time Creator Event
10:00 IG Started glass RPC 7up. Adjusted the current limits for the rest.
16:00 All Higher currents due to high temperature in the tent - The heating in the tent switched off

23.03.2010 - Tuesday

Time Creator Event
17:10 IG+SR Spotted increased humidity after the MS. Still at 5000 (compared to 40000 before the filter)

22.03.2010 - Monday

Time Creator Event
10:02 SR Reset of the GIF PC
  SR+IG The humidity over the weekend was >55%, the humidifier is empty, but we will not refill it, since the humidity is fine - 44%
22:30 RG gap 2up was in trip. HV restarted

20.03.2010 - Saturday

Time Creator Event
13:45 RG Reset of the GIF PC

19.03.2010 - Friday

Time Creator Event
09:30 IG 7up is again in over current (the limit was 100% over the current value). Left off for the weekend
    Glass RPC uGC/MS analysis done
  IG The C2H2F4 bootle was changed on the following dates:
    30/10/2009
    17/11/2009
    24/11/2009
    08/12/2009
    26/01/2010
    01/02/2010
    16/02/2010
    01/03/2010
    16/03/2010

18.03.2010 - Thursday

Time Creator Event
10:30 SR Reset of the GIF PC
  SR+IG The tent humidity is 58%. The humidifier to be set on lowest settingf, and if the RH is still to high, to be switched off completely.
  IG ATLAS HV channel ATL_NewSTD on at 9000 V & 10 uA
11:10 IG BigRPC2 channel showed almost no current at 5000 V, while BigRPC1 was showing ~40 uA. Switched off until Stefano finds out what is the problem. Meanwhile gap1/gap2/gap3 switched on together with Stefano - all looks fine for now.
14:30 IG Trip of 7up at 20 uA. Switched on and observed. In case new trip appears will be temporally switched off.
15:00 IG 7up current drop from 12 to 5 uA with no obvious reason
16:10 IG CAEN software reset
17:10 IG + SR MDT DAQ - stop, GIF - off, HV lowering humidifier refilling and setting to the lowest possible setting, GIF - on, HV - back to nominal, MDT DAQ - start
17:30 IG + SR Restart of GIF PC

17.03.2010 - Wednsday

Time Creator Event
12:00 SR Refilled the humidifier
18:20 IG Trip of 7up

16.03.2010 - Tuesday

Time Creator Event
09:00 All Problem with High voltage power supply. Error "No more channel in the group". In fact, I saw the same error on Sunday evening, but I didn't pay attention since it was the first time
11:00 All Only first module ON
14:00 All All our double gaps ON
18:00 RG since today there is some connection problem from GIF to the pc used for "daq". However, I checked directly on the pc and all the data are stored. I will investigate the problem tomorrow.
19:00 RG Also glass RPC ON

14.03.2010 - Saturday

Time Creator Event
21:30 RG High voltage restarted.

13.03.2010 - Saturday

Time Creator Event
11:30 RG gas system stopped during night most probably due to IR interlock. System restarted.

12.03.2010 - Friday

Time Creator Event
14:30 SR Refilled the humidifier, found the black one again in a state of error.

10.03.2010 - Wednesday

Time Creator Event
12:00 SR Refilled the humidifiers and lowered the set point of the one on top of the chambers a bit.
15:00 RG High voltage scan with ABS1

09.03.2010 - Tuesday

Time Creator Event
11:50 SR,RG High voltage scan with ABS5
14:37 SR,RG High voltage scan with ABS2
17:34 SR High voltage scan with source off
20:26 RG ABS1

08.03.2010 - Monday

Time Creator Event
8:30 SR Found the GIF PC on but not logged in.

05.03.2010 - Friday

Time Creator Event
7:20 IG System was off since 5:00. There was air comming in the open mode system (-1.2 mbar). System restarted, BINOS adjusted with -0.2. GC started on OMRETURN continuous analysis to see when the air returns back to normal levels.
    Test cartridge - taken for regeneration.
9:15 IG GIF PC restarted

04.03.2010 - Thursday

Time Creator Event
10:07 SR The humidity in the tent is 55%. The humidifiers should be readjusted.
    The air in the OMRETURN is 20 times less then yesterday, but still twice then the nominal value from Friday.
12:00 IG & SR Changed the settings of the humidifier. The black one was again not working when we found it.
15:30 - 18:30 IG+SR+RG+RF Dose map made of our RPCs to calculate the shadow that the MDT leaves on our chambers.
18:30 IG MDT DAQ - restarted, HV - nominal values (9250,9100), Glass RPC - off

03.03.2010 - Wednesday

Time Creator Event
16:00 RG from the data of this morning it seems that the air contamination in OMRET is decreasing much slower with respect to the fresh mixture (especially if we consider the 1 vol/h).
17:00 SR,RG new analysis show an air concentration of about 1/4 with respect this morning.
17:00 SR,RG the current with source off is about normal for DG in closed loop (1, 3, 6) and about double for DG in open mode (2, 4, 5)
17:30 SR,RG increased to +0.5 mbar the set-point for the open mode return (trying to avoid risk of underpressure in the chambers).
19:20 RG Restart of the GIF control PC

02.03.2010 - Tuesday

Time Creator Event
8:00 IG The system was in stop (since 23:00, yesterday) - the MFC was showing 0 for the C2H2F4 and normal values for the other components. Checking with the GC. The concentration of the nominal gas mixture components is the same as yesterday.
9:49 SR+IG Air 400 times then normal in OMRETURN, but currents are ok (?) - no change in CLRETURN
11:40 IG The total flow reduced with 3 l/h from the MFC, because might be that we are building up pressure between FIC82 and FIC21 - let's see..
19:12 IG Restart of the GIF control PC

01.03.2010 - Monday

Time Creator Event
7:21 IG 2dw on trip- restarting. All the rest - currents < 78 - going to 9250 V (9100 V for ch6).
    Installing the test cartridge again.
    The humidity in the tent went up to 60% on Saturday. Obviously the new configuration of the humidifiers results in even higher humidities. Going back to the old one (the humidifier with electronic setting over the CMS chambers, and the other on the chair)
8:02 IG Stefano requested short entrance. I used that to refill the humidifiers. The black one was still full, with error for missing water. Restarted the black one, refilled the other and left to minimum humidity setting, but on maximum on the top position.
    MDT DAQ restarted, chambers back to 9250 (9000) V. Currents < 80 (86 uA)
9:00 IG Air in the system found. The air comes from the supply of C2H2F4. Andre Castel promised to re-purge the supply bottle
15:30 IG Soft reset of the CAEN.
16:00 IG The air level is twice low - obviously the purging was successful.

27.02.2010 - Sunday

Time Creator Event
14:00 RG Since yesterday evening after re-start, the single gap corresponding to HV channel 7up shows high current. Switched off now. (How high? IG)
  RG Further investigation is needed. Tomorrow or better on Tuesday, we have to check gas and HV connections, when the source will be off for some time (I don't think a 5 minute access will be enough to understand the problem).
20:40 RG Also 7dw switched off.

26.02.2010 - Friday

Time Creator Event
10:40 SR Found the PC 30 pressure level at 1.2 again.
11:35 IG & SR System shutdown due to wrong gas mixture.
14:38 IG & SR System shutdown due to wrong gas mixture. There was no flow of freon through the Mass flow controller
    Problem was solved by increasing the presure of the Freon to 1 bar.
15:20 Gif shutdown to refill the humidifiers.
15:30 RG Freon supply pressure set to 0.7-0.8 bar.
19:00 SR RG HV back on

25.02.2010 - Thursday

Time Creator Event
17:50 IG High currents. Adjusted the voltage or current thresholds.

24.02.2010 - Wednsday

Time Creator Event
9:30 IG HV@9250V, currents < 90 uA, ch6@9050, currents < 87 uA
    Standby test cartridge found with only 100 mbar - refilled to 350 mbar.
    GIF control PC messages error sent today at 5:45. The PC was working fine.
  RG There was a network problem...in fact, the script was not able to retreive the data.
17:51 IG Removed Italian language from pccmsfwd09.cern.ch. Was the default language for log in/password typing, which was creating problems. Lemo cable on place.
    First results on nominal gas mixture (NGM) as standby gas - inconclusive: high leak in the cartridge + infrastructure, small amount of (possibly) polluted gas in the cartridge, long sampling line to the MS/GC results in no difference after the cartridge and the OMRETURN mixture. I'll prepare a custom, bug cartridge filled with MS4A and regenerate it tomorrow.

23.02.2010 - Tuesday

Time Creator Event
11:05 IG Start test of stanby phase with nominal gas mixture for MS4A. 400 mbar initial pressure
14:00 RG HV-I scan without radiation
15:30 IG Reduced humidifier and heater setting in the tent
18:45 RG radiation on; abs1

22.02.2010 - Monday

Time Creator Event
9:30 IG&SR Refilled the humidifiers and restarted the MDT data taking
21:00 RG problem with iC4H10: concentration down to 0. Kill send to HV system. I'm going to check.
21:20 RG normal gas mixture restored
  RG most probably the adjustment on the humidifiers have affected the mixer pressure
  RG iC4H10 pressure increased to 1 barg. Tomnorrow we will check the rotameters
22:15 RG HV: DG 1-5 at 9.2 kV; DG 6 at 9.0 kV

21.02.2010 - Sunday

Time Creator Event
13:30 RG HV: DG 1-5 at 9.2 kV; DG 6 at 9.0 kV

20.02.2010 - Saturday

Time Creator Event
21:30 RG Found HV off and gas system in stop.
    gas restarted. HV @ 6 kV untill tomorrow morning.

19.02.2010 - Friday

Time Creator Event
8:30 RG HV monitor pc and weather station pc restarted (most probably there was a nice update)
  RG HV restarted. For the moment 6 kV for all gaps
9:00 SR Pico crashed had to restart that computer
9:30 IG Restarting of the Pico
    Agreed on checking for CMF status regularly, in order to avoid loosing of data because of CMF reboots
    Since the HV is not ON, the GC analysis are left for the afternoon
10:00 IG HV@8800 V (found @8600 by Roberto) after increase currents < 45 uA
10:30 IG HV@9000 V after increase currents < 62 uA
11:20 IG HV@9200 V after increase currents < 90 uA
11:45 RG Final setting: 9200 V for gaps 1-5; 9000 V for gap 6
  RG Patm is very low, so for today we stay at lower HV

18.02.2010 - Thursday

Time Creator Event
9:10 IG&SR GIF communication pc restarted
  IG When restarting the PC, please wait for Windows to shut down. Sometimes, it is not able to stop the PVSS, and waits for acknowledge to stop the application.
9:30 IG&SR Source stopped to refill the humidifiers and replace the tape on the tent.
10:00 IG The currents are much lower today (atmospheric pressure back to 945 mbar). The current status is:
    HV for all is 9250V, Channel 6 - 9050V. Currents in the range of 54-88 uA (2dw - 14.9)
16:00 RG HV for 1-5 is 9200 V, Channel 6 - 9050V

17.02.2010 - Wednesday

Time Creator Event
8:30 SR The PPU & MS bakeout procedure stopped.
  SR GIF communication PC restarted
9.30 RG Channel 5+6 9050V; I about 92uA
10:15 IG GIF communication PC restarted
10.30 RG Patm 16/02 8.00 947 mbar; 17/02 10.30 939 mbar --> HV17 = 1.009 HV16; about 80-90 V higher
    conclusion: the decrease in Patm may justify the increase of current.
11:00 SR Start noticing a peak shift in the CLRETURN column C.
  IG It is possible to have peak shift after the bakeout of the columns yesterday. If the same peaks in a given column are all shifted (i.e. given component stays at the same retention times) - it is ok.
15:45 IG Overall currents continue to rise - 3dw in over current - limit change to 100 uA
    Changing all limits from 85/85/85/30/95/95/95/95/80/70/100/100 to 95/90/100/30/110/100/100/100/80/80/100/100
16:51 IG Changing voltage to 9200 and setting limits 85/85/90/20/100/95/95/95/70/70/100/100

16.02.2010 - Tuesday

Time Creator Event
9:10 IG The flow to the closed loop reduced to 0.5 l/h
9:40 RG Source off
10:00 IG The flow in HFC3 should always be 1/5 l/h
11:49 IG Start bakeout for all columns
11:57 IG Humidifier setting decreased to 40%, since the humidity in the tent is currently 55%
11:49 IG Start bakeout for all columns
13:42 RG Source on abs5
13:49 IG Finish bakeout for OV1. Continuing bakeout for all other columns
14:00 IG GIF communication PC reset by Roberto
16:45 RG Source abs2
19:45 RG Source abs1
20:00 IG After restarting the system with normal conditions, the currents in all the chambers are higher - not clear why - the temperature, humidity and pressure (945 -> 943) are the same.

15.02.2010 - Monday

Time Creator Event
09:35 IG GIF PC restarted
10:10 IG If the values in the MS for H2O, N2 and O2 are always zero, one should reload the default method and make Full tune of the device. This works. I have no idea if it is the right thing to do
14:00 IG Please find below a summary of the very short today GIF run meeting. Present: Bernhard, Sander, Roberto.
    *) General:
  Tomorrow (16/02) we will perform a scan with several absorption factor (1,2,5, radiation off). The time schedule of each absorption factor change will be agreed between MDT and RPC(Roberto).
  *) Set-up under test:
  ATLAS-MDT-RPC: Test on going. Contact person at cern: Stefanie Adomeit.
  ATLAS-RPC:
  Leak problem in the set-up still to be investigated (we can provide some help next week).
  CMS-RPC:
  Archana reported that there is no activity on the set-up. A pause is needed to perform some data analysis. Can we close the gas flow to your chambers?
  Gas group RPCs:
  Normal operation (both open and closed loop).
  Lyon RPC:
  Normal operation (open mode).
16:00 RG +SR humidifiers refilled

14.02.2010 - Sunday

Time Creator Event
21:00 IG GIF PC restarted

13.02.2010 - Saturday

Time Creator Event
0:15 RG GIF PC restarted
7:00 IG GIF PC restarted
12:00 RG GIF PC restarted

12.02.2010 - Friday

Time Creator Event
9:30 IG The GIF communication PC restarted. It was in a state in which the PC was up and running, but the default user (gifuser) was not logged in. Richard is investigating what was the reason for that.
    MS was showing 0 concentration on O2, N2, H2O - changing the method to a backup copy of the default method and doing tuning of the device help solving the problem - no idea why the problem appeared on the first place.
    The RH for both OM and CL is ~35%. I increased it a little bit.
    HV: Channels 1-5 at the nominal 9250V all currents are in the range of 46-73 uA (2dw - 11.7 uA) Channel 6 @ 9100V @ 79 uA
10:57 SR Pressure at the pc30 dropped to -0.4 open the valve to get it to 0.3 again.
18:40 IG Again repeating environmental data from the GIF - I restarted the PC

11.02.2010 - Thursday

Time Creator Event
9:30 RG GIF communication PC restarted
  IG from Richard: This time it was a PVSS problem
9:54 IG GIF restart for humidifier refill
    Lower humidifier set to lower value of humidity
    MDT data taking restart
    HV: Channels 1-5 at 9200V all currents are in the range of 51-78 uA (2dw - 13 uA) Channel 6 @ 9050V @ 83 uA

10.02.2010 - Wednesday

Time Creator Event
9:30 IG&SR Reduced the total flow in the OM to have 6 l/h for all chambers (was 7.5 l/h)
    The air in the OMRETURN is lower then yesterday, but still 10% more then the levels before Christmas
    The HV for channels 1-5 raised to 9150 V resulting in currents in the range of 47-72uA (not considering channel 2 dw, the current in which is as usual lower)
    RH ~ 35% - ok. In the tent is 45% - the setting in the humidifier should be changed on the next access. (Or maybe switch the places of the humidifiers? In the moment the humidifier with automatic switch-off is on the top of Stefano's cart)
    Tunning of the MS done - the counter restarted.
  RG The presence of air is due to the under-pressure in the open mode return. Air is entering the system through the double way bubbler. Try to stabilize the pressure and to minimize the change in the purifiers rack.
  IG No Roberto, that is not the case. The pressure was negative only for short time, and did not result in significant increase of the air in the system. In fact we know that low pressure at PC30 means air in the system, and we are monitoring it closely.
  IG Good to know (based on Roberto's e-mail):
    The minimum value of PC30 (the pressure on the chambers in open mode) is -1.6 mbar, after which the bubbler xBUB21 starts bubbling air in the chambers and the purifiers in open mode. The pressure at the chambers in this case is -2.0 mbar. The only way toaffect this prtessureis by adjusting PI41, PCV12, HV13 or open mode pump speed.
11:37 SR The computer with the soprane and MSD software rebooted itself.
18:34 IG GIF communication PC restarted - the logging of the environmental data works again

09.02.2010 - Tuesday

Time Creator Event
09:30 IG Stop MS for pump oil exchange.
  IG The web site for our pump: http://www.pfeiffer-vacuum.com/cnt/en/368/
  IG The operating instructions: https://twiki.cern.ch/twiki/pub/LHCgas/GCMS/Pump_Operating_Instructions.pdf
  IG The next oil + filter exchange should be done on 09.08.2010
15:00 RG Set max HV for gap 11 and 12 (DG 6) to 9.0 kV. Currents are still too high
16:20 SR Found the PC30 pressure level at -1.5.
17:09 SR Soprane processing software crashed restarded the MSD software and Soprane again

08.02.2010 - Monday

Time Creator Event
09:00 IG&SR Pico had crashed - restarted
  IG&SR RH both in open and in CL - low (~29%) - adjusted
  IG&SR MS message "Limits fail: Pump oil days" - The oil of the low pressure pump should be exchanged by the end of the day.
09:30 IG All double gaps voltage to 8000 V
10:35 IG&SR Short stop of GIF.
  IG MDT data taking restarted.
  IG&SR Refilling of both humidifiers with demineralized water.
  IG&SR Starting the first barrel of demineralized water.
  IG Plastic valve for the water barrel taken from the workshop.
11:30 SR High amount of air in the OMRETURN - Probably due to the system down over the weekend. ( 3 times higher then the SF6 in PPU column)
14:00 SR Low pressure on PC30 (-1.6) - due to change in setting of PI41
14:39 SR & IG RH still low (< 32 %) readjusted.

07.02.2010 - Sunday

Time Creator Event
20:00 RG System was in stop - restarted

06.02.2010 - Saturday

Time Creator Event
10:00 RG The loop - closed

05.02.2010 - Friday

Time Creator Event
16:00 IG The demineralized water moved in the storage area. (transport request was needed)
  IG Huge amount of argon the CLRETURN - due to argon in the humidifier not flushed in the open mode. Flushing at least until tomorrow.

04.02.2010 - Thursday

Time Creator Event
10:00 IG Conditioning of the filters for the closed loop:
  IG MS - (after 22 hours) Air going down still (MS), still higher then the SF6 (PPU). Impurities - constant (1,3,7,8)
  IG R11 - (after 22 hours) Air - ok (OV1)
12:00 IG Argon found in the CL (~10%). Opening the loop from HV90
18:09 IG Restarting normal operation with purifiers in closed loop - not stable yet - the humidity for OM/CL and the pressure in the OM. Not all chambers fully ON.
19:00 RG The chambers' voltages lowered to 6000 V - the currents are too high

18.01.2010 - Monday

Time Creator Event
15:30 IG Summary from the GIF meeting
  *) General:
  We decided to dedicate one day every week (Tuesday) to maintenance and calibration. In detail, every Tuesday we will go through the following steps:
  From 8.30 am to 2.00 pm Source OFF (access to GIF)
  From 2.00 pm to 4.00 pm radiation ON (ABS 5)
  From 4.00 pm to 6.00 pm radiation ON (ABS 2)
  From 6.00 pm back to normal operation: radiation ON (ABS 1)
  During next week, we will see if it is really needed one day every week or may be it is enough every two weeks.
  *) Present configuration:
  Run with Argon
  No flammable gas available
  Today we did the safety visit with Glimos and FGSO. Tomorrow we will complete the safety inspection and, hopefully, we will be able to restart operation with the standard RPC mixture very soon.
  *) Set-up under test:
  ATLAS-MDT-RPC: Test re-started. Test will continue al least until the end of February
  ATLAS-RPC: For economical reasons, it will be nice to test the closed loop circulation and fix possible problem in the present configuration (Ar flow). In any case, the output of the input safety bubbler should be connected to the exhaust.
  CMS-RPC: No news. Stefano (Colafranceschi), please let us know if you have particular requirements.
  Gas group RPCs: Run with Argon. Ready to restart normal operation both in closed and open mode.
  Lyon RPC: HV scan will be performed with longer time interval between each step.

Dates and Maps

MS Status
Time Creator Event
18.10.2010 SR MS 5A+4A installed in the closed loop.
29.07.2010 SR MS4A+5A installed in the loop.
17.05.2010 IG MS 5A+4A changed with "old==standard" cartridge
11.05.2010 All MS 5A+4A changed with a temporary small cartridge (until the "old" cartridge is in regeneration)
07.05.2010 All MS saturated 5A+4A
31.03.2010 All MS 5A+3A saturated and changed
06.02.2010 All MS started

Cu Status
Time Creator Event
06.02.2010 All Cu started

Ni Status
Time Creator Event
06.02.2010 All Ni started

15.03.2010 - IG&RC - remove NiAl from the closed loop

09.08.2010 - IG - The MS pump oil and filter should be exchanged

Map of gas and high voltage connections RG 23/5/2010
DG HVMod HVCable Gas Open/Closed
0 A6 A8 Rack1/FIC13 Open
1 A1 A2 1/11 Closed
2 A2 A3 4/403 Open
3 A3 A4 1/12 Closed
4 A4 A5 4/402 Open
5 A5 A6 4/401 Open
Glass A7 A7 Gas Open

-- IvanGlushkov - 08-Feb-2010

Edit | Attach | Watch | Print version | History: r235 < r234 < r233 < r232 < r231 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r235 - 2011-07-08 - BeatriceMandelli
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LHCgas All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright &© 2008-2023 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback